Suppose the demand curve is linear and is given by the equation P = a – bQ where P is price and Q is quantity. What is the consumer surplus if the equilibrium price is P* and equilibrium quantity is Q*?


1
Expert's answer
2021-07-29T03:05:02-0400

Consumer Surplus = Total Utility – (Price x Quantity)


"Demand is given by : p = a - bQ"


"Marginal revenue(MR) = \frac{d(TR)}{dQ}"


"where TR = p\u00d7Q = (a - bQ)\u00d7Q"


"= aQ - bQ^2"


"Thus, MR = \frac{d(TR)}{dQ} = a - 2bQ"


"Hence, MR = a - 2bQ"


"When Q = 0 ,"


"then, P = a - b\u00d70 = a"


"MR = a - 2b\u00d70 = a"


Consumer surplus;

"aQ\u2212bQ ^\n2 - (a \u00d7 0)"


"= aQ\u2212bQ ^\n2"


= 0-0 = 0